Operations Report – May 23rd

Crew 299 Operations Report 5-23-2024

SOL: 11

Name of person filing report: Rishabh Pandey (Engineer)

Non-Nominal Systems: Suit 9 (broken visor), Suit 11 (fan issue)

Notes on non-nominal systems: Issues with both suits were noted before our mission, both will not be used during our mission.

Rovers Used: Curiosity
Start hours:183.8
Start charge:100
End charge: 76
Currently charging: yes

Start hours:190.5
Start charge:100
End charge: 80
Currently charging: yes

General notes on rovers: None

Summary of Hab operations:
      WATER USE: 19 Gallons
      Water (static tank): 218.3 (using effective value from spreadsheet)
      Static tank pipe heater: off
      Static tank heater: off
      Toilet tank emptied: no

Summary of internet usage: Catch up on Earthly news, email responses, code development, entertainment

Summary of suits and radios: Radios 1-6 used. Suits 7,8,and 10 used

Summary of GreenHab operations: NA, GreenHab out of operation

Summary of ScienceDome operations: More rock samples were made and analyzed. Radishes were watered and checked in on. Soil was further analyzed.

Summary of RAM operations: Measuring tape used.
Summary of observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site